OpenStudy (anonymous):

(systems) what is the solution to the following system? y=1/2x+3 x=15-y

OpenStudy (anonymous):

Re-writing it \[y=\frac{ 1 }{ 2 }x+3\] \[x=15-y\]

OpenStudy (anonymous):

You can use substitution, from the first equation y is equal to (1/2)x+3 so you can take that and plug it into the second equation where y is x=15-(1/2)x+3 now solve for x

OpenStudy (anonymous):

So all i do is solve for x and y and then im done?

OpenStudy (chihiroasleaf):

yes.., to find the solution, you need to find the x and y ... you have y = (1/2) x + 3 and x = 15 - y if you want to use substitution method, substitute the first equation into the second equation so, you'll have x = 15 - { (1/2)x + 3} what is x?

OpenStudy (anonymous):

im a little stuck. I distributed the negative so i got x=15-1/2x-3 then i continue and get x=12-1/2x i have i done it right so far? and how do i continue?

OpenStudy (nottim):

what you need how to continue?

OpenStudy (nottim):

gotta be more sure of yourself, thats for sure

OpenStudy (anonymous):

@NotTim what do you mean? uh.. i kinda need some help on this were not all geniuses

OpenStudy (nottim):

well...add likes to likes really.